OpenStudy (anonymous):

If y(x) = 4x, what is x when y(x) = 4? A. 1 B. 4 C. 2 D. 0

OpenStudy (anonymous):

so first set up the problem. 4(x)=4

OpenStudy (anonymous):

then divide 4 from both sides 4/4(x)=4/4 1(x)=1 or x=1

OpenStudy (anonymous):

hope that helped :-)